My son needs help. A coffee merchant has coffee beans that sell at $9 per pound and $12 per pound. The two types are to be mmixed to create 100lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture
Answer by stanbon(75887)

My son needs help. A coffee merchant has coffee beans that sell at $9 per pound and $12 per pound. The two types are to be mmixed to create 100lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture
-------------
$9 coffee DATA:
Let amount be "x" lbs ; Value of this is 9x dollars
--------------
$12 coffee DATA:
Amount = "100-x" lbs ; Value of this is 12(100-x) = 1200-12x dollars
---------------
Mixture DATA:
Amount = 100 lbs ; Value of this is 11.25*100 = 1125 dollars
---------------
EQUATION:
value + value = 1125
9x + 1200-12x = 1125
-3x = -75
x = 25 lbs. ( amt of $9 coffee in the mixture)
100-x = 75 lbs ( amt of $12 coffee in the mixture)
===============
Cheers,
Stan H.